# Fractional Chern Insulator and Quantum Anomalous Hall Crystal in Twisted MoTe2

### v2

@@ Abstract
Recent experimental advances have uncovered fractional Chern insulator (FCI) states in twisted MoTe2 (tMoTe2) systems under zero magnetic field. Understanding the interaction effects on topological phases within realistic model presents a significant theoretical challenge. Here, we construct a moiré lattice model tailored for tMoTe2 and conduct investigations using state-of-the-art tensor-network methods. Our ground-state calculations reveal a rich variety of interaction-driven and filling-dependent topological phases, including FCIs, Chern insulators, and generalized Wigner crystals, which were found in recent experiments. For FCI state, dynamical simulations uncover a single-particle excitation continuum with a finite charge gap, reflecting the fractionalized charge excitations. Finite-temperature calculations further determine characteristic charge activation and ferromagnetic transition temperatures, reconciling existing experimental discrepancies. Furthermore, using a realistic moiré lattice model, we predict the emergence of quantum anomalous Hall crystals exhibiting integer Hall conductivity at fractional fillings in tMoTe2. By integrating ground-state, finite-temperature, and dynamical analyses, our work establishes a comprehensive framework for understanding correlated topological phases in tMoTe2 and related moiré systems.
@@ I Introduction
Twisted bilayer transition metal dichalcogenides, such as twisted MoTe2 (tMoTe2), have attracted significant research interest due to the emergence of correlated topological states [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. The observation of fractional Chern insulator (FCI)  [1, 2, 3, 4, 5] represents a hallmark that has long been sought in condensed matter physics. Notably, FCI states at filling factors $\nu_{h}=-2/3$, $-3/5$, $-4/7$, $-4/9$, and $-5/9$ have been experimentally observed recently in tMoTe2 at zero magnetic field [7, 8, 9, 10, 11, 12]. It not only advances the understandings of correlated topological quantum matter, but also shed light on fault-tolerant topological quantum computation. Recent experiments have revealed signatures of (fractional) excitations in FCIs, including conducting edge states [13], charge gaps [12], and thermal activation gaps [10]. Consequently, the exotic quantum states in tMoTe2 — particularly correlated topological phases such as FCI and their generalizations — have attracted significant theoretical attention. [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@
While charge orders typically compete with FCIs, we highlight that their intricate interplay can also stabilize novel quantum phases of matter. Prominent examples include quantum Hall crystals (QHCs) in strong magnetic fields [33, 34, 35, 36, 37] and, more remarkably, quantum anomalous Hall crystals (QAHCs) that emerge even at zero field [38]. Both states exhibit an integer quantized Hall conductivity despite a fractional electron filling. QAHCs further distinguish themselves by simultaneously breaking time-reversal, space group, and lattice translation symmetries. While examples of QHCs and QAHCs have been observed in multilayer graphene systems [39, 40, 41, 42, 43], the microscopic mechanisms that stabilize these exotic states – such as electron correlations and symmetry-breaking patterns – are under active investigations [44, 24, 32, 45].
@@
In particular, the exploration of QAHCs in semiconductor moiré systems has garnered considerable attention, with theoretical proposals predicting their emergence in tMoTe2 [24] and AB-stacked MoTe2/WSe2 heterobilayers [32]. Nevertheless, elucidating these correlated topological states and studying their phase transitions presents substantial theoretical difficulties, owing to their highly entangled many-body character. Exact diagonalization (ED) are limited to small clusters ($\sim$30 sites) [24], and Hartree-Fock (HF) methods fail to capture entangled phases like FCIs [32]. More importantly, although a wide range of experiment measurements have been performed, including spectroscopy [8, 13, 11, 46], transport [9, 10], thermodynamics [7], magnetometry [12], and microstructural characterization [47], the present theoretical efforts have mainly focused on the ground-state properties, leaving a critical gap between theory and experiments [44].
To address these fundamental questions and bridge the existing gap, we first construct a realistic Hubbard-type model for tMoTe2 in real space (see Fig. 1), constructed from properly chosen Wannier orbitals. This real-space approach provides a distinct perspective from prevailing numerical studies of moiré systems, which predominantly employ momentum-space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 20] or hybrid real-space/momentum-space representations [48, 49, 50, 31]. We then perform comprehensive, unbiased large-scale tensor-network calculations to study experimentally relevant properties of tMoTe2, including ground-state, finite-temperature, and dynamical characteristics.
In the ground-state study, we reveal the FM order and experimental discovered FCIs for various electron fillings $11/3\leq\nu<$ 1, or hole fillings $-1<\nu_{h}\leq-1/3$ via particle-hole transformation (as explained below), alongside a competition between CO and FCI states (see Fig. 1e). Remarkably, QAHC states are revealed at $\nu=1/2,2/3,3/5,$ and $4/5$, arising from two distinct mechanisms — hole crystal formation and band folding. We further propose that these QAHC states could be transitioned into FCI states by melting their charge order through tuning parameters like dielectric constant and temperature.
Our finite-temperature calculations further reveal a non-monotonic behavior of FM transition temperature $T_{c}$ versus fillings, in agreement with experimental observations [8]. Furthermore, we also compute the single-particle spectral function at zero temperature, for several representative fillings including $\nu=1/2$, $2/3$, and $1$, and identify the charge gaps as approximately 18.2 meV, 8.4 meV, and 6.2 meV, respectively. Finite-temperature simulations further yield results of the local density of state at the Fermi level $A_{\rm loc}(\omega=0)$, which disclose a typical thermal activation temperature of about $T^{*}\approx 30$ K for $\nu=2/3,1$. Notably, the zero-temperature charge gaps are found to be significantly larger than the energy scales of thermal activation and FM transition, a phenomenon that has been also observed in experiments but lacks theoretical understanding [12, 10]. Our numerical results reproduce these experimental observations and resolve their contradictions with the realistic real-space model, which establish a robust foundation for further theoretical and experimental investigations on MoTe2 and moiré materials in general.
II Model and Methods
II.1 Real-Space Model on the Honeycomb Lattice
@@
Figure 1: Real-space model on the honeycomb lattice and its global phase diagram. a Schematic representation of the model with hopping parameters ($t_{1}$, $t_{2}$, $t_{3}$) and interactions ($U$, $V_{1}$, $V_{2}$, $V_{3}$) up to the third nearest neighbor on a honeycomb lattice. $\bm{a}_{1}=(\sqrt{3}/2,1/2)$ and $\bm{a}_{2}=(0,1)$ denote the primitive vectors of the honeycomb lattice. Numerical simulations employ a X-cylinder (XC) geometry with periodic boundary conditions along $\bm{a}_{2}$ direction. Panels b and c show real-space electronic density profile of exponentially localized Wannier orbitals, normalized to the peak density. $\rho^{(l)}_{i\uparrow}=\big{|}w^{(l)}_{i\uparrow}\big{|}^{2}$ for spin-$\uparrow$ is centered on site $i=(A,B)$ in layer $l=(b,t)$. The Wannier orbitals reside mostly at the honeycomb lattice sites, corresponding to the XM/MX stacking regions of the twisted bilayer MoTe2. d Magnetic phase diagram in the $\epsilon_{r}$-$\nu$ plane, where the dashed line (obtained with $N_{y}=5$ cylinder) and dash-dotted boundary lines ($N_{y}=4$ data) separate the FMz and non-FMz phases. e Phase diagrams with representative fillings ($\nu=1,2/3,1/2$, and $1/3$) contains various phases, including the CI, FCI, QAHC, and CO phases. The phase diagram is constructed using width-6 ($N_{y}=6$) simulations, a system size that accommodates the root configurations of both FCIs and composite Fermi liquids [51, 52].
@@
Bilayer MoTe2 with a twist angle of approximately 3.89 has recently attracted significant experimental interest [7, 10, 9, 8, 12, 11]. To provide a theoretical understanding, density functional theory (DFT) calculations [6, 53, 18, 19, 54] have been conducted to reveal the relevant low-energy degrees of freedom. They comprise two bands per valley with Chern numbers +1 and -1, respectively. Due to the spin-valley locking [55], the spin-$\uparrow$ state is associated with valley $\mathbf{K}$, while the spin-$\downarrow$ state is linked to valley $\mathbf{K}^{\prime}$, resulting in a minimal model encompassing four electronic bands. For simplicity, we refer primarily to the spin states throughout the manuscript, omitting explicit reference to the valley degree of freedom. It has been shown that the low energy physics of tMoTe2 can be effectively described by a continuum model [6] (see also SM I [56]), which can be mapped to a tight-binding model through a proper choice of Wannier basis, through the trial state procedure [57, 29, 23]. As illustrated in Fig. 1b and c, the Wannier orbitals $w^{(l)}_{i\uparrow}$ for spin-$\uparrow$ are exponentially localized on the two sublattices ($i=A$ or $B$) of a honeycomb lattice [57] and exhibit clear layer resolution ($l=b$ or $t$, meaning bottom or top). Specifically, $w^{(b)}_{A\uparrow}$ predominantly resides on the bottom layer, while $w^{(t)}_{B\uparrow}$ resides on the top layer. The situation is the same for spin-$\downarrow$ orbitals since the two spin orientations are related by the time-reversal symmetry.
@@
By projecting the kinetic term and Coulomb interactions onto these Wannier orbitals, we derive a Hubbard-type real-space model for tMoTe2 on the honeycomb lattice, as depicted in Fig. 1a. The Hamiltonian reads
@@
where $c^{\dagger}_{i\sigma}$ ($c_{i\sigma}$) creates (annihilates) an electron with spin $\sigma=\uparrow$, $\downarrow$ on site $i$. The density at site $i$ is $n_{i}=n_{i\uparrow}+n_{i\downarrow}$, with $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}$. $\mu$ is the chemical potential, and the hopping parameter $t_{ij}$ is between the $(i,j)$ pair for electrons with spin-$\uparrow$, conjugate to those for spin-$\downarrow$. We consider the hopping amplitudes ($t_{1},t_{2},t_{3}$), up to the third nearest neighbor in our model (see Fig. 1a). Notably, the next-nearest neighbor hopping $t_{2}$ for spin-$\uparrow$ electrons along the direction depicted in Fig. 1a incorporates a phase factor of $\phi=2\pi/3$. Therefore, hopping parameters (for spin-$\uparrow$) are $t_{1}=-3.225$ meV, $t_{2}=2.120e^{i2\pi/3}$ meV, and $t_{3}=0.947$ meV.
@@
The onsite Coulomb interaction is denoted as $U=1192.71/\epsilon_{r}$ meV, where $\epsilon_{r}$ represents the relative dielectric constant (see SM I [56]). For the further couplings $V_{ij}$, we consider a dual-gate screened Coulomb interaction in tMoTe2, i.e.,
$V(r)=\frac{e^{2}}{4\pi r\epsilon_{0}\epsilon_{r}}e^{-r/\xi}=\frac{276.92}{\epsilon_{r}r/L_{m}}\exp\Big(-\frac{r/L_{m}}{\xi/L_{m}}\Big{missing})~{}\text{meV},$ (2)
@@
where $\epsilon_{0}$ is the vacuum permittivity, $r$ is the distance between two sites, and $\xi$ is the screening length. $L_{m}\simeq 5.2$ nm represents the length of primitive vectors of the moiré lattice (see Fig. 1a), which is taken as the unit of length in the present study. In our simulations, we mainly consider $10\leq\epsilon_{r}\leq 20$ in tMoTe2, aligning with previous studies [18, 26, 23]. Additionally, we adopt a reasonable value of $\xi/L_{m}=2$, as also used in Ref. [23], and retain interaction terms up to the third nearest neighbor [i.e., $V_{1},V_{2},V_{3}$ as indicated in Fig. 1a], compatible with our finite width of cylinder (up to 6). Such a truncation in $V_{ij}$ (and also $t_{ij}$) turns out to be a good approximation, even for large $\xi=30~{}{\rm nm}\simeq 5.77~{}L_{m}$, as illustrated in SM I [56].
Since the onsite Coulomb interaction $U$ exceeds the hopping parameters $t_{1},t_{2},t_{3}$ by over an order of magnitude, we can project out double occupancy at low temperatures. The numerical benchmarks between two models are shown in the SM III [56]. Then we arrive at a spinful $t$-$V$ model whose Hamiltonian reads
${}\begin{split}H=&\sum_{i<j}\big{(}t_{ij}P_{i}c^{\dagger}_{i\uparrow}c_{j\uparrow}P_{j}+t^{*}_{ij}P_{i}c^{\dagger}_{i\downarrow}c_{j,\downarrow}P_{j}+h.c.\big{)}\\ -&\mu\sum_{i}n_{i}P_{i}+\sum_{i<j}V_{ij}P_{i}n_{i}n_{j}P_{j},\end{split}$ (3)
In the calculations of fully spin-polarized cases, e.g., FCI and QAHC states, we further simplify the model to a spinless version. This is justified as Eq. [3] conserves electron numbers for both spin flavors, and the spinless Hamiltonian is given by
${}\begin{split}H=&\sum_{i<j}\big{(}t_{ij}c^{\dagger}_{i}c_{j}+h.c.\big{)}-\mu\sum_{i}n_{i}+\sum_{i<j}V_{ij}n_{i}n_{j},\end{split}$ (4)
where $c^{\dagger}_{i}$ creates an electron at site $i$, and $n_{i}=c^{\dagger}_{i}c_{i}$ is the particle number operator. This spinless $t$-$V$ model is employed in the calculation of Hall conductivity, and its validity is confirmed by comparing the results with those from the Hubbard-type model [1] or the spinful $t$-$V$ model [3].The three models and their usage in various simulations are summarized in Table 1.
Model No. $d$ $T=0$ $T>0$ Dynamics
Hubbard-type Eq. [1] 4 $\times$
spinful $t$-$V$ Eq. [3] 3 $\times$
spinless $t$-$V$ Eq. [4] 2 $\times$
@@
Here for the sake of simplicity, we have performed a particle-hole transformation in the model [1]. It inverts the hopping parameters and chemical potential, thus the non-interacting band structure [19] (see also SM I [56]). In the following, we use $\nu$ to denote the electron filling per unit cell, which consists of two lattice sites, $A$ and $B$. Consequently, the experimentally relevant hole filling $-1\leq\nu_{h}\leq 0$ thus corresponds to the electron filling $0\leq\nu\leq 1$ in our practical model calculations. Moreover, to reduce computational costs, the Hubbard-type model described by Eq. [1] can be simplified to either a spinful or spinless $t$-$V$ model (see Materials and Methods).
II.2 Tensor Network Methods
In this study, we investigate the zero-temperature, finite-temperature, and dynamical properties using state-of-the-art tensor-network methods on an X-cylindrical (XC) geometry of size $N_{y}\times N_{x}\times 2$.
After obtaining the ground state, we employ real-time evolution with the TDVP approach [62, 63] to compute the retarded Green’s function and subsequently derive the single-particle spectral function via Fourier transformation. In the computation of dynamical properties, it is necessary to determine the chemical potential $\mu$ beforehand in order to fix the position of the Fermi level $(\omega=0)$. In the present study, we estimate $\mu$ based on low-temperature calculations (practically at $T\simeq 1.45$K), utilizing the $\mu$-$\nu$ relation (see, e.g., Fig.6f). In practice, a Parzen window is applied to suppress non-physical oscillations [64, 65], achieving an energy resolution of $2.67$ meV (with evolution time up to 3 meV-1). We choose time step $\delta t=1/16$ meV-1 and bond dimension $\chi=1000$, which renders accurate results in the TDVP calculations. The data convergence and more technical details are presented in the SM IV [56].
The tanTRG method [66] is employed to study the finite-temperature properties of the system. An accurate matrix product operator representation of the density matrix, with bond dimension $\chi$ up to 2000, enables us to investigate systems with sizes up to XC4×24×242424$\times$ 24$\times$ 24 × 24 × 2 and down to a low temperature of $T\simeq 1.45$ K (i.e., $T/t_{1}\simeq 0.0375$). In our tanTRG calculations, we employ the grand canonical ensemble and implement a specially designed chemical potential $\mu$-tuning technique to precisely control the particle number. To study the FM transition, we introduce a small pinning field $h=0.01$ to stabilize the FM order in the finite-size systems. To improve the efficiency, we implement Abelian spin and/or charge symmetries in the simulations and adopt controlled bond expansion techniques [67, 68] to speed up the finite-temperature calculations.

### v3

@@ Abstract
Recent experimental advances have uncovered fractional Chern insulator (FCI) states in twisted MoTe2 (tMoTe2) systems under zero magnetic field. Understanding the interaction effects on topological phases within realistic model presents a significant theoretical challenge. Here, we construct a moiré superlattice model tailored for tMoTe2 and conduct investigations using state-of-the-art tensor-network methods. Our ground-state calculations reveal a rich variety of interaction-driven and filling-dependent topological phases, including FCIs, Chern insulators, and generalized Wigner crystals, which are revealed in recent experiments. For FCI state, dynamical simulations uncover a single-particle excitation continuum with a finite charge gap, reflecting the fractionalized charge excitations. Finite-temperature calculations further determine characteristic charge activation and ferromagnetic transition temperatures, reconciling existing experimental discrepancies. Furthermore, using this realistic lattice model, we predict the presence of quantum anomalous Hall crystals exhibiting integer Hall conductivity at fractional fillings in tMoTe2. By integrating ground-state, finite-temperature, and dynamical analyses, our work establishes a comprehensive framework for understanding correlated topological phases in tMoTe2 and related moiré systems.
@@ I Introduction
Twisted bilayer transition metal dichalcogenides, such as twisted MoTe2 (tMoTe2), have attracted significant research interest due to the emergence of correlated topological states [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. The observation of fractional Chern insulator (FCI)  [1, 2, 3, 4, 5] represents a hallmark that has long been sought in condensed matter physics. Notably, FCI states at filling factors $\nu_{\mathrm{h}}=-2/3$, 3/5-3/5, 4/7-4/7, 4/9-4/9, and 5/9-5/9 have been experimentally observed recently in tMoTe2 at zero magnetic field [7, 8, 9, 10, 11, 12]. It not only advances the understandings of correlated topological quantum matter, but also shed light on fault-tolerant topological quantum computation [14]. Recent experiments have revealed signatures of (fractional) excitations in FCIs, including conducting edge states [13], charge gaps [12], and thermal activation gaps [10]. Consequently, the exotic quantum states in tMoTe2 — particularly correlated topological phases such as FCI and their generalizations — have attracted significant theoretical attention  [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@
While charge orders typically compete with FCIs, we highlight that their intricate interplay can also stabilize novel quantum phases of matter. Prominent examples include quantum Hall crystals (QHCs) in strong magnetic fields [34, 35, 36, 37, 38] and, more remarkably, quantum anomalous Hall crystals (QAHCs) that emerge even at zero field [39]. Both states exhibit an integer quantized Hall conductivity despite a fractional electron filling. QAHCs further distinguish themselves by simultaneously breaking time-reversal, space group, and lattice translation symmetries. While examples of QHCs and QAHCs have been observed in multilayer graphene systems [40, 41, 42, 43, 44], the microscopic mechanisms that stabilize these exotic states – such as electron correlations and symmetry-breaking patterns – are under active investigations [45, 25, 33, 46]. Indeed, the term “QAHC” is used here in a generalized sense, following its broader usage in recent studies of moiré systems. This usage differs from the original definition, which referred to spontaneous crystalline quantum anomalous Hall states in systems without an underlying periodic potential.
@@
In particular, the exploration of QAHCs in semiconductor moiré systems has garnered considerable attention, with theoretical proposals predicting their emergence in tMoTe2 [25] and AB-stacked MoTe2/WSe2 heterobilayers [33]. Nevertheless, elucidating these correlated topological states and studying their phase transitions presents substantial theoretical difficulties, owing to their highly entangled many-body character. Exact diagonalization (ED) are limited to small clusters ($\sim$30 sites) [25], and Hartree-Fock methods fail to capture highly entangled phases like FCIs [33]. More importantly, although a wide range of experiment measurements have been performed, including spectroscopy [8, 13, 11, 47], transport [9, 10], thermodynamics [7], magnetometry [12], and microstructural characterization [48], prior theoretical efforts have mainly focused on the ground-state properties, leaving a critical gap between theory and experiments [45].
@@
Figure 1: Real-space model on the honeycomb lattice and its global phase diagram. Panels (a) and (b) show real-space electronic density profile of exponentially localized Wannier orbitals, normalized to the peak density. $\rho^{(l)}_{i\uparrow}=\big|w^{(l)}_{i\uparrow}\big|^{2}$ for spin-$\uparrow$ is centered on site $i=(A,B)$ in layer $l=(b,t)$. The Wannier orbitals reside mostly at the honeycomb lattice sites, corresponding to the XM/MX stacking regions of the twisted bilayer MoTe2. (c) Schematic representation of the model with hopping parameters ($t_{1}$, $t_{2}$, $t_{3}$) and interactions ($U$, $V_{1}$, $V_{2}$, $V_{3}$) up to the third nearest neighbor on a honeycomb lattice. $\bm{a}_{1}=(\sqrt{3}/2,1/2)$ and $\bm{a}_{2}=(0,1)$ denote the primitive vectors of the honeycomb lattice. Numerical simulations employ a X-cylinder (XC) geometry with periodic boundary conditions along $\bm{a}_{2}$ direction. (d) Magnetic phase diagram in the $\epsilon_{\mathrm{r}}$-$\nu$ plane, where the dashed line (obtained with $N_{y}=5$ cylinder) and dash-dotted boundary lines ($N_{y}=4$ data) separate the FMz and non-FMz phases. (e) Phase diagrams with representative fillings ($\nu=1,2/3,1/2$, and $1/3$) contains various phases, including the CI, FCI, QAHC, and CO phases. The phase diagram is constructed using width-6 ($N_{y}=6$) simulations, a system size that accommodates the root configurations of both FCIs and composite Fermi liquids [49, 50].
To address these fundamental questions and bridge the existing gap, we first construct a realistic Hubbard-type model for tMoTe2 in real space (see Fig. 1), constructed from properly chosen Wannier orbitals. This real-space approach provides a distinct perspective from prevailing numerical studies of moiré systems, which predominantly employ momentum-space [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] or hybrid real-space/momentum-space representations [51, 52, 53, 32]. We then employ comprehensive, unbiased, large-scale tensor-network calculations to investigate the ground-state, finite-temperature, and dynamical properties, connecting our results with experimental findings on tMoTe2.
In our ground-state analysis, we identify the emergence of ferromagnetic (FM) order and FCI states across electron fillings $1/3\leq\nu<1$, in agreement with experimental observations, while also revealing a competitive interplay between CO and FCI phases. Through a particle-hole transformation — detailed further below — this filling range corresponds to hole fillings of $1<\nu_{h}\leq-1/3$. Remarkably, QAHC states are revealed at $\nu=1/2,2/3,3/5,4/5$ and $0.63$, arising from two distinct mechanisms — hole crystal formation and band folding. We further propose that these QAHC states could be driven into FCI states by melting their charge order through tuning parameters like dielectric constant and temperature.
Our finite-temperature calculations further reveal a non-monotonic behavior of FM transition temperature $T_{c}$ versus fillings, in agreement with experimental observations [8]. Furthermore, we also compute the single-particle spectral function at zero temperature, for several representative fillings including $\nu=1/2$, $2/3$, and $1$, and identify the charge gaps as approximately 18.2, 8.4, and 6.2 meV, respectively. Finite-temperature simulations further yield results of the local density of state at the Fermi level $A_{\rm loc}(\omega=0)$, which disclose a typical thermal activation temperature of about $T^{*}\approx 30$ K for $\nu=2/3,1$. Notably, the zero-temperature charge gaps are found to be significantly larger than the energy scales of thermal activation and FM transition, a phenomenon that has been also observed in experiments but lacks theoretical understanding [12, 10]. Our numerical results reproduce these experimental observations and resolve their contradictions with a realistic real-space model, which establish a robust foundation for further theoretical and experimental investigations on MoTe2 and moiré materials in general.
II Real-space model on the honeycomb lattice
@@
Bilayer MoTe2 with a twist angle of approximately 3.89 has recently attracted great research interest [7, 10, 9, 8, 12, 11]. To provide a theoretical understanding, density functional theory (DFT) calculations [6, 54, 19, 20, 55] have been conducted to reveal the relevant low-energy degrees of freedom. They comprise two bands per valley with Chern numbers +1+1 and 1-1, respectively. Due to the spin-valley locking [56], the spin-$\uparrow$ state is associated with valley $\mathbf{K}$, while the spin-$\downarrow$ state is linked to valley $\mathbf{K}^{\prime}$, resulting in a minimal model encompassing four electronic bands. For simplicity, we refer primarily to the spin states throughout the manuscript, omitting explicit reference to the valley degree of freedom. It has been shown that the low energy physics of tMoTe2 can be effectively described by a continuum model [6] (see also Supplementary section I online), which can be mapped to a tight-binding model through a proper choice of Wannier basis, through the trial state procedure [57, 30, 24]. As illustrated in Figs. 1a and b, the Wannier orbitals $w^{(l)}_{i\uparrow}$ for spin-$\uparrow$ are exponentially localized on the two sublattices ($i=A$ or $B$) of a honeycomb lattice [57] and exhibit clear layer resolution ($l=b$ or $t$, meaning bottom or top). Specifically, $w^{(b)}_{A\uparrow}$ predominantly resides on the bottom layer, while $w^{(t)}_{B\uparrow}$ resides on the top layer. The situation is the same for spin-$\downarrow$ orbitals since the two spin orientations are related by the time-reversal symmetry.
@@
By projecting the kinetic term and Coulomb interactions onto these Wannier orbitals, we derive a Hubbard-type real-space model for tMoTe2 on the honeycomb lattice, as depicted in Fig. 1c. The Hamiltonian reads
@@
where $c^{\dagger}_{i\sigma}$ ($c_{i\sigma}$) creates (annihilates) an electron with spin $\sigma=\uparrow$, $\downarrow$ on site $i$. The density at site $i$ is $n_{i}=n_{i\uparrow}+n_{i\downarrow}$, with $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}$, and $\mu$ is the chemical potential. The hopping $t_{ij}$ is defined for each $(i,j)$ bond in the spin-$\uparrow$ sector, with the corresponding terms for spin-$\downarrow$ obtained by Hermitian conjugation. We include hopping amplitudes up to the third nearest neighbor ($t_{1},t_{2},t_{3}$) in our model (see Fig. 1c). Notably, the next-nearest neighbor hopping $t_{2}$ for spin-$\uparrow$ electrons along the direction depicted in Fig. 1c incorporates a phase factor of $\phi=2\pi/3$. Therefore, hopping parameters (for spin-$\uparrow$) are $t_{1}=-3.225$ meV, $t_{2}=2.120e^{i2\pi/3}$ meV, and $t_{3}=0.947$ meV.
@@
The onsite Coulomb interaction is denoted as $U=1192.71/\epsilon_{\mathrm{r}}$ meV, where $\epsilon_{\mathrm{r}}$ represents the relative dielectric constant (see Supplementary section I online). For further couplings $V_{ij}$, we consider a dual-gate screened Coulomb interaction in tMoTe2, i.e.,
$V(r)=\frac{e^{2}}{4\pi r\epsilon_{0}\epsilon_{\mathrm{r}}}e^{-r/\xi}=\frac{276.92}{\epsilon_{\mathrm{r}}r/L_{m}}\exp\Big(-\frac{r/L_{m}}{\xi/L_{m}}\Big.)~\text{meV},$ (2)
@@
where $\epsilon_{0}$ is the vacuum permittivity, $r$ is the distance between the two sites (i,j), and $\xi$ is the screening length. $L_{m}\simeq 5.2$ nm represents the length of primitive vectors of the moiré superlattice (see Fig. 1c), which is taken as the unit of length in the present study. In our simulations, we mainly consider $10\leq\epsilon_{\mathrm{r}}\leq 20$ in tMoTe2, aligning with previous studies [19, 27, 24]. In practice, we adopt a reasonable value of $\xi/L_{m}=2$, as also used in Ref. [24], and retain interaction terms up to the third nearest neighbor (i.e., $V_{1},V_{2},V_{3}$ as indicated in Fig. 1c), compatible with our finite width of cylinder (up to 6). Despite the larger experimental $\xi\simeq 30~{\rm nm}(\simeq 5.77L_{m})$, this truncation in $V_{ij}$ and $t_{ij}$ remains a valid approximation, supported by analysis in Supplementary Section I.
@@
In this study, we investigate the zero-temperature [58, 59, 60, 61], finite-temperature [62], and dynamical [63, 64, 65, 66, 67, 68] properties using state-of-the-art tensor-network methods (see Supplementary section II online). The simulations are conducted on an X-cylindrical (XC) geometry of size $N_{y}\times N_{x}\times 2$. Here for the sake of simplicity, we have performed a particle-hole transformation in the model (1). It inverts the hopping parameters, chemical potential, and thus the non-interacting band structure [20] (see also Supplementary section I online). In the following, we use $\nu$ to denote the electron filling per unit cell, which consists of two lattice sites, $A$ and $B$. Consequently, the experimentally relevant hole filling $1\leq\nu_{h}\leq 0$ thus corresponds to the electron filling $0\leq\nu\leq 1$ in our practical model calculations. Moreover, to reduce computational costs, the Hubbard-type model described by Eq. (1) can be simplified to either a spinful or spinless $t$-$V$ model (Supplementary section II online).